Net Present Value Method, Internal Rate of Return Method, and Analysis

The management of Quest Media Inc. is considering two capital investment projects. The estimated net cash flows from each project are as follows:

Year Radio Station TV Station
1 $380,000 $720,000
2 380,000 720,000
3 380,000 720,000
4 380,000 720,000
Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.352 2.991
6 4.917 4.355 4.111 3.784 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

The radio station requires an investment of $983,820, while the TV station requires an investment of $2,055,600. No residual value is expected from either project.

Required:

1a. Compute the net present value for each project. Use a rate of 10% and the present value of an annuity of $1 in the table above. If required, use the minus sign to indicate a negative net present value. If required, round to the nearest whole dollar.

Radio Station TV Station
Present value of annual net cash flows $ $
Less amount to be invested $ $
Net present value $ $

1b. Compute a present value index for each project. If required, round your answers to two decimal places.

Present Value Index
Radio Station
TV Station

2. Determine the internal rate of return for each project by (a) computing a present value factor for an annuity of $1 and (b) using the present value of an annuity of $1 in the table above. If required, round your present value factor answers to three decimal places and internal rate of return to the nearest whole percent.

Radio Station TV Station
Present value factor for an annuity of $1
Internal rate of return % %

3. The net present value, present value index, and internal rate of return all indicate that the radio station is a better financial opportunity compared to the tv station , although both investments meet the minimum return criterion of 10%.
Answer #1
1a
Radio Station TV Station
Present value of annual net cash flows 1204600 2282400
Less amount to be invested 983820 2055600
Net present value 220780 226800
1b
Present Value Index
Radio Station 1.22 =1204600/983820
TV Station 1.11 =2282400/2055600
2
Radio Station TV Station
Present value factor for an annuity of $1 2.589 2.855
Internal rate of return 20% 15%
3
The net present value, present value index, and internal rate of return all indicate that the radio station is a better financial opportunity compared to the tv station , although both investments meet the minimum return criterion of 10%.
Workings:
Radio Station TV Station
Present value of annual net cash flows =380000*3.170 =720000*3.170
Radio Station TV Station
Present value factor for an annuity of $1 =983820/380000 =2055600/720000